OpenStudy (anonymous):

Determine if the following infinite series is convergent or divergent:

OpenStudy (anonymous):

\[\sum_{n=2}^{\infty} 1/\ln n^lnlnn\]

OpenStudy (anonymous):

hint:ln x is less than sqrt x

OpenStudy (amoodarya):

use integral test hint use u=ln n

OpenStudy (anonymous):

is it \(\frac{1}{\ln(n)\ln(\ln(n))}\)?

OpenStudy (anonymous):

\[(\ln)^\ln(\ln n)=e^\ln(\ln n)^\ln(\ln n)=e^[\ln(\ln n)]^2\]

OpenStudy (anonymous):

and by the hint ln(ln n) is less than sqrt(ln n) implies....1/e[ln(ln n)]squared is greater than 1/e^ln n all equals to 1/n

OpenStudy (anonymous):

i can't figure out what it is you are writing, but it diverges for sure, because the log grows very slowly

OpenStudy (anonymous):

it diverges and im sorry for the dashes in my work i meant to put n in those so sorry for the sloppiness

OpenStudy (anonymous):

maybe \[\frac{1}{\ln(n)^{\ln(ln(n))}}\]?

OpenStudy (anonymous):

correct

OpenStudy (amoodarya):

 it is converegent because